# Principled RLHF from Heterogeneous Feedback via Personalization and Preference Aggregation

### v1

@@ 4.1 Algorithms
Algorithm 3 outputs $K$ clustered policies and a map from human users to clusters. The input of the algorithm is $\widehat{\mathcal{D}}=\cup_{i\in[N]}\widehat{\mathcal{D}}_{i}$ where $\widehat{\mathcal{D}}_{i}=\{(o_{i}^{(j)},\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})_{j\in[N_{p}]}\}$, which is the same as Algorithm 1. After estimating the representation parameter $\widehat{\omega}$, the algorithm will estimate the reward function parameters $(\widehat{\theta}_{(k)})_{k\in[K]}$ with Equation 4.2. Lastly, we find the best policy based on expected value function.
We also provide a practical algorithm that uses DPO and EM (Moon, 1996) algorithms to solve Equation 4.2. Given the inherent complexity of this hierarchical optimization problem, which presents more challenges than standard optimization tasks (Anandalingam and Friesz, 1992), we propose a novel algorithm that circumvents the need for explicit reward function estimation in Algorithm 4. Our approach begins by randomly assigning each human user to a cluster. Subsequently, we reassign random human users to the cluster where the policy most effectively maximizes their empirical DPO loss (Equation 4.3). Finally, we refine our solution by optimizing the DPO loss function for the selected human users within each cluster, thereby enhancing the overall policy effectiveness.
@@
Algorithm 4 ClusterDPO: Learning $K$ clustered policies by DPO
  Randomly initialize $f^{0}(i)$ for $i\notin\{p_{1},\dots,p_{K}\}$
  for $t\in[T]$ do
     Randomly select $K$ human users $p_{1},\dots,p_{K}$.
     for $i\in[N]$ do
        if $i\notin\{p_{1},\dots,p_{K}\}$ then
           Define $f^{t}(i)\leftarrow f^{t-1}(i)$
        end if
     end for
     Assign $f^{t}(p_{k})$ for all $k\in[K]$ as
$f^{t}(p_{k})\leftarrow\underset{s\in[K]}{\arg\max}\sum_{j\in[N_{p}]}\log\sigma\left(\beta\log\frac{\pi_{(s)}^{t-1}(a_{p_{k},0}^{(j)}\mid s_{p_{k}}^{(j)})}{\pi^{\text{old}}(a_{p_{k},0}^{(j)}\mid s_{p_{k}}^{(j)})}-\beta\log\frac{\pi_{(s)}^{t-1}(a_{p_{k},1}^{(j)}\mid s_{p_{k}}^{(j)})}{\pi^{\text{old}}(a_{p_{k},1}^{(j)}\mid s_{p_{k}}^{(j)})}\right)$ (4.3)
and obtain $\pi_{(s)}^{t}$ for all $s\in[K]$.
  end for
  Assign $f^{T+1}(i)$ for all $i\in[N]$ as
  Output: $({\pi}_{(k)}^{T})_{k\in[K]}$ and $f^{T+1}$
4.2 Results and Analyses
First, we introduce label discrepancy (Mohri and Muñoz Medina, 2012) for the preference dataset and reward function class. We defer all the proofs of this section to Appendix D.
@@ 4.2 Results and Analyses
(Total Expected Value Function Gap). Suppose Assumptions 3.1, 3.2, 3.3, and 3.4 hold. Also, assume that $C_{r}(\mathcal{G}_{r},\pi,\mu_{i,\text{ref}},i)\leq C_{\text{max}}^{\prime}$ for all policy $\pi$ and $i\in[N]$. For any $\delta\in(0,1]$, all $i\in[N]$ and $\lambda>0$, with probability at least $1-\delta$, the output $((\widehat{\pi}_{(k)})_{k\in[K]},\widehat{f})$ of Algorithm 3 satisfies
@@
$\leq cN\kappa\Biggl{(}{\frac{\log(2K/\delta)}{N_{p}}}+{\frac{kK\log(N_{p}/K)}{N_{p}}}+{\frac{k\xi^{2}\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{r}}(1/(NN_{p}))/\delta)}{{NN_{p}}}}$
$\qquad\qquad\qquad+\left(\sum_{i\in[N]}\frac{1}{N}\texttt{disc}(\mathcal{D}_{i},\mathcal{C}_{\widehat{f}(i)},\mathcal{G}_{\psi^{\star}}))\right)^{2}+\left(\frac{\log(\mathcal{N}_{\mathcal{G}_{\psi^{\star}}}(1/NN_{p})/\delta)}{NN_{p}}\right)^{2}\Biggr{)}^{1/4},$
@@
We note that due to the $\sqrt{dk/N_{p}}$ order on the right-hand side of Lemma 1, we have a slower rate in Theorem 4.1 than Theorem 3.4. This gap is mainly due to the fact that the analysis of Lemma 1 should cover uniformly for arbitrary $\widehat{f}$, and also due to a difference between $\max$ and expectation of $\max$, which is bounded using McDiarmid’s inequality.
@@
In contrast to the results in Section 3, we additionally assume $C_{r}(\mathcal{G}_{r},\pi,\mu_{1},i)\leq C_{\text{max}}^{\prime}$ in Theorem 4.1. To adopt a pessimistic approach, constructing a confidence set for clustered reward functions across all clusters is necessary. However, the ambiguity of which human user belongs to which cluster complicates this analysis, as pessimism would need to be applied to every potential cluster. Consequently, defining a confidence set for every possible clustering scenario is required, significantly complicating the analysis of the algorithm.
@@ 5 Reward and Preference Aggregation
This section adheres to the RLHF setting with a single LLM, while handling the heterogeneous human feedback by reward/preference aggregation. For reward aggregation, we leverage comparison data similar to the methods discussed in the previous sections. We first estimate individual reward functions based on such data and then aggregate these functions to form a unified reward model. In comparison, for preference aggregation, we introduce a novel framework termed “probabilistic opinion pooling”. Specifically, instead of relying on binary comparison data, human users provide feedback as probability vectors. This approach eliminates the need to aggregate heterogeneous preferences via reward functions, allowing for the direct aggregation of probabilistic opinions provided by users.
5.1 Reward Aggregation with Comparison Data
We introduce the following reward aggregation rules (Equations (5.3) and (5.6)), which are favorable as they satisfy several pivotal axioms in social choice theory. These axioms – monotonicity, symmetry, continuity, independence of unconcerned agents, translation independence, and the Pigou-Dalton transfer principle – are crucial for ensuring fairness and consistency in the decision-making process (List, 2013; Skiadas, 2009, 2016). We present the definition of these axioms in Section E.1 for completeness. The aggregation rules are presented as follows:
@@ 5.1 Reward Aggregation with Comparison Data
where $\bm{r}=(r_{1},\dots,r_{N})^{\intercal}$ is a reward vector with trajectory input. Note that Equation 5.3 and Equation 5.6 are equivalent in the sense of the associated optimal policy, as the $\log$ function is monotonically increasing. We can verify that $\lim_{\alpha\to-\infty}\text{Agg}_{\alpha}(\bm{r})=\min_{i\in[N]}r_{i}$ and $\lim_{\alpha\to\infty}\text{Agg}_{\alpha}(\bm{r})=\max_{i\in[N]}r_{i}$. This implies that when $\alpha$ is small, the reward aggregation rule emphasizes on $\min_{i\in[N]}r_{i}$, and when $\alpha$ is large, it emphasizes on $\max_{i\in[N]}r_{i}$. When $\alpha=0$, Equation 5.3 represents utilitarianism, and when $\alpha\to-\infty$, Equation 5.3 represents a Leximin-based aggregation rule (List, 2013).
@@ 5.1.1 Algorithm and Analysis
(Expected Value Function Gap). Suppose Assumptions 3.1, 3.2, 3.3, and 3.4 hold. For any $\delta\in(0,1]$, all $i\in[N]$ and $\lambda>0$, with probability at least $1-\delta$, the output $\widehat{\pi}$ of Algorithm 5 satisfies
@@
where $c_{\alpha}>0$ is a constant depending on $\alpha$, and other constants are defined in Section 3.1.
We defer the proof of Theorem 5.1 to Section E.2. Lastly, we consider the tightness of our analysis by providing a theoretical lower bound of the sub-optimality gap of aggregation.
@@
${\rm CB}(\Lambda)\coloneqq\left\{Q\coloneqq\left(\left\{\mu_{0},\mu_{1}\right\},\{\tau_{i,0}^{(j)},\tau_{i,1}^{(j)}\}_{i\in[N],j\in[N_{p}]},\omega,\boldsymbol{\theta}\right){\,\big{|}\,}C_{\bm{r}}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\Lambda\text{ for all }i\in[N]\right\}$
is the family of MDP with $N$ reward functions and $H=1$ instances.
@@
(Zhong et al., 2024) also considered reward aggregation rules, adhering to the axiom of scale independence but not translation independence (Moulin, 2004). Here, scale/translation independence of the aggregation rule means that the aggregation rule should yield the same choice even if the reward functions are scaled/translated, respectively. Our theoretical results can also be extended to the reward aggregation rule in (Zhong et al., 2024). However, since we also consider the relationship of the reward aggregation rule and preference aggregation rule by probabilistic opinion pooling, which will be presented in the next section, we only present Equation 5.3 for the reward aggregation rule.
(Zhong et al., 2024) also considered Nash’s bargaining (Nash, 1953), which maximizes $\prod_{i\in[N]}(r_{i}-\min r_{i})$ rather than $\prod_{i\in[N]}r_{i}$. In this case, they can also consider translation independence. Ours can also be extended by substituting $r_{i}$ to $r_{i}/\min r_{i}$ in the aggregation function (Equation 5.3), but we decided not to include it in our paper since we did not have this result in our initial draft.
@@ 5.2.1 RLHF with Probabilistic Opinion Feedback
Interestingly, Equation 5.10, which describes the aggregation of probabilistic preferences, has a connection to Equation 5.6, concerning reward aggregation, under the assumption of the Plackett-Luce model for the relationship between reward functions and preference models (Definition 2.1). We then formalize the connection between the probabilistic opinion pooling in Equation 5.10 and the reward aggregation rule in Equation 5.3. We defer the proof of Theorem 5.3 to Section E.4.
@@ 5.2.2 Algorithm
Now, we provide an algorithm that uses the feedback in the form of probabilistic opinions (Algorithm 6). The only difference from the DPO algorithm (Rafailov et al., 2024) is to change the deterministic answer $a_{i}$ to the $a_{i}$ sampled based on the probabilistic opinion pooling, which is in the second line in the for loop of Algorithm 6.
@@
5.3 Mechanism Design for Preference Aggregation
@@ 5.3.1 Setup
In this setup, we will first prove the existence of a cost function $c_{i}:\Delta(\mathcal{A})^{N}\to\mathbb{R}$ for all $i\in[N]$ that induces truthful reporting of probabilistic opinions from human labelers. Here, the input of $c_{i}$ is the probabilistic opinion of every human labeler. This is also called the dominant strategy incentive-compatible (DSIC) mechanism (Nisan and Ronen, 1999; Börgers, 2015; Roughgarden, 2010). Then, we prove that there exists an aggregation rule and cost function that induce DSIC, and also maximize social welfare. We denote each human labeler’s underlying (true) probabilistic opinion as $p_{i}\left(s^{\left(j\right)}\right)$ for each question $s^{(j)}$. Accounting for such cost, we define the utility function of individual $i$ for question $s^{(j)}$ as
Here, $d:\Delta(\mathcal{A})\times\Delta(\mathcal{A})\to\mathbb{R}$ represents the distance between the underlying true probabilistic opinion and the aggregated preference. Moreover, we define the welfare function of individual $i$ from addressing question $s^{(j)}$ as $\text{Wel}_{i}^{(j)}(O)=-d(p_{i}(s^{(j)}),O)$ for any $O\in\Delta(\mathcal{A})$.
@@
We can instantiate $d(p,q)$ as the KL-divergence. Also, we may instantiate $d_{\alpha}(p,q)=\text{sgn}(\alpha)\frac{1}{1-\alpha}\sum_{j\in\mathcal{A}}\left(1-p_{j}^{\alpha}q_{j}^{1-\alpha}\right)$, which is a variant of the $\alpha$-Renyi divergence for $\alpha\neq 0$. One can easily check that $d_{\alpha}(p,q)\geq 0$. In fact, one can also prove that $\lim_{\alpha\to 1}d_{\alpha}(p,q)=d(p,q)$ with $d(p,q)$ being the KL-divergence (Section E.5).
5.3.2 Mechanism and Guarantees
@@
Unfortunately, the classical VCG mechanism presents certain limitations such as it cannot be solved in polynomial time in general (Nisan and Ronen, 1999; Börgers, 2015; Roughgarden, 2010). We here adopt certain forms of allocation rule (which corresponds to the aggregation rule in our RLHF setting) and cost functions as follows, which allow the outcome set to be a simplex (with infinitely many outcomes):
@@
If we assume the relationship between reward and preference follows the PL model (Definition 2.1), then Equation 5.3 implies a welfare-maximizing aggregation rule, which connects reward aggregation and mechanism design. We defer all proofs for the results in Section 5.3.2 to Section E.6.

### v2

@@ 3.2.1 Algorithms
Algorithm 3 outputs $K$ clustered policies and a map from human users to clusters. The input of the algorithm is $\widehat{\mathcal{D}}=\cup_{i\in[N]}\widehat{\mathcal{D}}_{i}$ where $\widehat{\mathcal{D}}_{i}=\{(o_{i}^{(j)},\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})_{j\in[N_{p}]}\}$, which is the same as Algorithm 1. After estimating the representation parameter $\widehat{\omega}$, the algorithm will estimate the reward function parameters $(\widehat{\theta}_{(k)})_{k\in[K]}$ with Equation 3.6. Lastly, we find the best policy based on the expected value function. We defer a practical algorithm that uses DPO (Rafailov et al., 2024) (and also refer to Appendix D) and EM (Moon, 1996) algorithms to solve Equation 3.6 to Algorithm 6.
@@
3.2.2 Results and Analyses
To analyze the clustering-based personalization approach, we adapt the notion of label discrepancy in (Mohri and Muñoz Medina, 2012) to our RLHF setting, for preference data and a given reward function class. We defer all the proofs of this section to Appendix G.
@@ 3.2.2 Results and Analyses
The discrepancy is defined as the supremum value of the difference between the log-likelihood of the preference data when taking expectations over two human dataset distributions. This quantity will be used in the analysis to characterize the gap between the log-likelihood of the estimated parameters and the underlying parameters. A similar concept is frequently used in domain adaptation (Mansour et al., 2009) and federated learning (Mansour et al., 2020).
@@
(Total Expected Value Function Gap). Suppose Assumptions 1, 2, 3, and 4 hold. Also, assume that $C_{r}(\mathcal{G}_{r},\pi,\mu_{i,\text{ref}},i)\leq C_{\text{max}}^{\prime}$ for all policy $\pi$ and $i\in[N]$. For any $\delta\in(0,1]$, all $i\in[N]$ and $\lambda>0$, with probability at least $1-\delta$, the output $((\widehat{\pi}_{(k)})_{k\in[K]},\widehat{f})$ of Algorithm 3 satisfies
@@
$\leq cN\kappa\Biggl{(}\underbrace{\frac{\log(2K/\delta)}{N_{p}}+{\frac{kK\log(N_{p}/k)}{N_{p}}}}_{(i)}+{\frac{k\xi^{2}\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{r}}(1/(NN_{p}))/\delta)}{{NN_{p}}}}$
$\qquad\qquad\qquad+\underbrace{\left(\sum_{i\in[N]}\frac{1}{N}\texttt{disc}(\mathcal{D}_{i},\mathcal{C}_{\widehat{f}(i)},\mathcal{G}_{\psi^{\star}}))\right)^{2}}_{(ii)}+\left(\frac{\log(\mathcal{N}_{\mathcal{G}_{\psi^{\star}}}(1/NN_{p})/\delta)}{NN_{p}}\right)^{2}\Biggr{)}^{1/4},$
@@
Note that Theorem 3.4 addresses the bias-variance tradeoff: as the number of clusters ($K$) increases, term (i) (variance) increases, while term (ii) (bias) decreases. Also, we note that due to the $\sqrt{kK/N_{p}}$ order on the right-hand side of Lemma 1, we have a slower rate in Theorem 3.4 than Theorem 3.1. This gap is mainly due to the fact that the analysis of Lemma 1 should cover uniformly for arbitrary $\widehat{f}$, and also due to a difference between $\max$ and expectation of $\max$, which is bounded using McDiarmid’s inequality.
@@
In contrast to the results in Section 3.1, we additionally assume $C_{r}(\mathcal{G}_{r},\pi,\mu_{1},i)\leq C_{\text{max}}^{\prime}$ in Theorem 3.4. To adopt a pessimistic approach, constructing a confidence set for clustered reward functions across all clusters is necessary. However, the ambiguity of which human user belongs to which cluster complicates this analysis, as pessimism would need to be applied to every potential cluster. Consequently, defining a confidence set for every possible clustering scenario is required, significantly complicating the analysis of the algorithm.
@@ 4 Reward and Preference Aggregation
This section adheres to the RLHF setting with a single LLM, while handling the heterogeneous human feedback by reward/preference aggregation. For reward aggregation, we first estimate individual reward functions and then aggregate these functions to form a unified reward model. In comparison, for preference aggregation, we introduce a novel framework termed “probabilistic opinion pooling”. Specifically, instead of relying on binary comparison data, human users provide feedback as probability vectors. This approach eliminates the need to aggregate heterogeneous preferences via reward functions, allowing for the direct aggregation of probabilistic opinions provided by users.
4.1 Reward Aggregation
We introduce the following reward aggregation rules (Equations (4.3) and (4.6)), which are favorable as they satisfy several pivotal axioms in social choice theory. These axioms – monotonicity, symmetry, continuity, independence of unconcerned agents, translation independence, and the Pigou-Dalton transfer principle – are crucial for ensuring fairness and consistency in the decision-making process (List, 2013; Skiadas, 2009, 2016). We present the definition of these axioms in Section H.1 for completeness. The aggregation rules are presented as follows:
@@ 4.1 Reward Aggregation
where $\bm{r}=(r_{1},\dots,r_{N})^{\intercal}$ is a reward vector with trajectory input. Note that Equation 4.3 and Equation 4.6 are equivalent in the sense of the associated optimal policy, as $\log(x)$ is monotonically increasing. We can verify that $\lim_{\alpha\to-\infty}\text{Agg}_{\alpha}(\bm{r})=\min_{i\in[N]}r_{i}$ and $\lim_{\alpha\to\infty}\text{Agg}_{\alpha}(\bm{r})=\max_{i\in[N]}r_{i}$. This implies that when $\alpha$ is small (or large), the reward aggregation rule emphasizes $\min_{i\in[N]}(\text{or }\max_{i\in[N]})r_{i}$, respectively. When $\alpha=0$, Equation 4.3 represents utilitarianism, and when $\alpha\to-\infty$, Equation 4.3 represents a Leximin-based aggregation rule (List, 2013).
@@ 4.1.1 Algorithm and Analysis
(Expected Value Function Gap). Suppose Assumptions 1, 2, 3, and 4 hold. For any $\delta\in(0,1]$, all $i\in[N]$ and $\lambda>0$, with probability at least $1-\delta$, the output $\widehat{\pi}$ of Algorithm 4 satisfies
@@
where $c_{\alpha}>0$ is a constant depending on $\alpha$, and other constants are defined in Section 3.1.1.
We defer the proof of Theorem 4.1 to Section H.2. Lastly, we prove the tightness of our analysis by providing a theoretical lower bound of the sub-optimality gap of aggregation.
@@
${\rm CB}(\Lambda)\coloneqq\left\{Q\coloneqq\left(\left\{\mu_{0},\mu_{1}\right\},\{\tau_{i,0}^{(j)},\tau_{i,1}^{(j)}\}_{i\in[N],j\in[N_{p}]},\omega,\boldsymbol{\theta}\right){\,\big{|}\,}C_{\bm{r}}^{\prime}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\Lambda\text{ for all }i\in[N]\right\}$
is the family of MDP with $N$ reward functions and $H=1$ instances. $C_{\bm{r}}^{\prime}$ is defined in Equation 3.4.
@@
(Zhong et al., 2024) also considered reward aggregation rules, adhering to the axiom of scale independence but not translation independence (Moulin, 2004). Here, scale/translation independence of the aggregation rule means that the aggregation rule should yield the same choice even if the reward functions are scaled/translated, respectively. Our theoretical results can also be extended to the reward aggregation rule in (Zhong et al., 2024). However, since we also consider the relationship of the reward aggregation rule and preference aggregation rule by probabilistic opinion pooling, which will be presented in the next section, we only present Equation 4.3 for the reward aggregation rule. (Zhong et al., 2024) also considered Nash’s bargaining (Nash, 1953), which maximizes $\prod_{i\in[N]}(r_{i}-\min r_{i})$ rather than $\prod_{i\in[N]}r_{i}$. In this case, they can also consider translation independence. Ours can also be extended by substituting $r_{i}$ to $r_{i}/\min r_{i}$ in the aggregation function (Equation 4.3), but we decided not to include it in our paper since we did not have this result in our initial draft.
@@ 4.2.1 RLHF with Probabilistic Opinion Feedback
Interestingly, Equation 4.10, which describes the aggregation of probabilistic preferences, has a connection to Equation 4.6, concerning reward aggregation, under the assumption of the Plackett-Luce model for the relationship between reward functions and preference models (Definition 4.1). We then formalize the connection between the probabilistic opinion pooling in Equation 4.10 and the reward aggregation rule in Equation 4.3. We defer the proof of Theorem 4.3 to Section H.4.
Definition 4.1.
@@ 4.2.2 Algorithm
We provide an algorithm that uses the feedback in the form of probabilistic opinions (Algorithm 5). The only difference from the DPO algorithm (Rafailov et al., 2024) is to change the deterministic answer $a_{i}$ to the $a_{i}$ sampled based on the probabilistic opinion pooling, which is in the second line in the for loop of Algorithm 5.
@@
5 Mechanism Design for Preference Aggregation
@@ 5.1 Setup
In this setup, we will first prove the existence of a cost function $c_{i}:\Delta(\mathcal{A})^{N}\to\mathbb{R}$ for all $i\in[N]$ that induces truthful reporting of probabilistic opinions from human labelers. Here, the input of $c_{i}$ is the probabilistic opinion of every human labeler. This is also called the dominant strategy incentive-compatible (DSIC) mechanism (Nisan and Ronen, 1999; Börgers, 2015; Roughgarden, 2010). Here incentive-compatibility means for labeler $i$, whatever the other labelers’ reports are, truthful reporting will always maximize her own utility function. Intuitively, our mechanism punishes labeler $i$ through cost $c_{i}$ for the externality she posed on other labelers, i.e., if she makes other labelers worse off based on the aggregation outcome.
Moreover, we prove that this cost function $c_{i}$ not only induce DSIC but also maximize social welfare. We denote each human labeler’s underlying (true) probabilistic opinion as $p_{i}\left(s^{\left(j\right)}\right)$ for each question $s^{(j)}$. Accounting for such cost, we define the quasi-linear utility function of individual $i$ for question $s^{(j)}$ as
Here, $d:\Delta(\mathcal{A})\times\Delta(\mathcal{A})\to\mathbb{R}$ represents the distance between the underlying true probabilistic opinion and the aggregated preference. Moreover, we define the welfare function of individual $i$ from addressing question $s^{(j)}$ as $\text{Wel}_{i}^{(j)}(O)=-d(p_{i}(s^{(j)}),O)$ for any $O\in\Delta(\mathcal{A})$. In other words, we define utility function as subtracting cost from the welfare of individual $i$. The concepts of utility and welfare discussed in this section should not be confused with the reward used in previous sections. Here, welfare pertains to the intrinsic rewards (preferences) of individuals and the aggregation of heterogeneous human rewards (preferences).
@@
We can instantiate $d(p,q)$ as the KL-divergence. Also, we may instantiate $d_{\alpha}(p,q)=\text{sgn}(\alpha)\frac{1}{1-\alpha}\sum_{j\in\mathcal{A}}\left(1-p_{j}^{\alpha}q_{j}^{1-\alpha}\right)$, which is a variant of the $\alpha$-Renyi divergence for $\alpha\neq 0$. One can easily check that $d_{\alpha}(p,q)\geq 0$. In fact, one can also prove that $\lim_{\alpha\to 1}d_{\alpha}(p,q)=d(p,q)$ with $d(p,q)$ being the KL-divergence (Section H.5).
5.2 Mechanism and Guarantees
@@
The private valuation $v_{i}$ is corresponding to our welfare function for $i$th individual. Unfortunately, the classical VCG mechanism presents certain limitations such as it cannot be solved in polynomial time in general (Nisan and Ronen, 1999; Börgers, 2015; Roughgarden, 2010). We here adopt certain forms of allocation rule (which corresponds to the aggregation rule in our RLHF setting) and cost functions as follows, which allow the outcome set to be a simplex (with infinitely many outcomes):
@@
If we assume the relationship between reward and preference follows the PL model (Definition 4.1), then Equation 4.3 implies a welfare-maximizing aggregation rule, which connects reward aggregation and mechanism design. We defer all proofs for the results in Section 5.2 to Section H.6.
Remark 6.
An analog of Theorem 5.1 and Theorem 5.2 can also be applied to reward aggregation. Additionally, under the PL model, the mechanism design for reward aggregation and preference aggregation coincide.
6 Experiments
We now conduct an empirical evaluation of our methods’ performance on a text summarization task, using the Reddit TL;DR summarization dataset and the Reddit TL;DR human feedback dataset (Stiennon et al., 2020). For the Reddit TL;DR summarization dataset, (Stiennon et al., 2020) filtered the TL;DR summarization dataset (Völske et al., 2017) to ensure quality. The Reddit TL;DR human feedback dataset is constructed with two components: comparison and axes evals. The comparison component contains labeled comparisons between pairs of summaries with workers identified by unique IDs, while the axes evals component contains ratings of summaries along three axes: accuracy, coverage, and coherence. We used GPT-J 6B (Wang and Komatsuzaki, 2021) and LLaMA3 8B models (Meta, 2024) in our experiments.
Figure 2: Accuracy of different methods with 3 times experiments. P: Personalized, C: Clustered, G(L): General (Linear) representation. Naive RLHF: original training method.
In Section 6.1, we fine-tuned the personalized reward model with Algorithm 1 and Algorithm 3, without pessimism. We ranked workers based on the number of annotated comparisons in the training split of the dataset and included the top 5 workers for training. To balance the number of samples for each worker, we took the worker with the fewest samples among the top 5 as the baseline. We then randomly sampled the same number of comparisons from the other workers so that each worker had 5,373 comparison samples, resulting in a total of 26,865 samples for training. Similarly, for the validation set, we applied the same method. We randomly sampled the same number of comparisons as the worker with the fewest samples from the top 5 workers used in training. Each worker had 1,238 samples for validation, resulting in a total of 6,190 samples for validation.
In Section 6.2, we fine-tuned the personalized reward model using Algorithm 4, without incorporating pessimism. We considered three types of reward functions: accuracy-reward, coverage-reward, and coherence-reward. Since this dataset is only publicly available for the validation set (with 8,585 samples) and the test set (with 6,313 samples), we used the validation set for fine-tuning the training set of our model and validated it with the samples in the test set. We defer other details in Appendix I.
6.1 Experiment 1: Reward Model Performance with comparison Dataset
In Experiment 1, we compared our Algorithm 1 and Algorithm 3 with naive RLHF methods. We constructed a reward model using a supervised fine-tuned language model and added a linear layer to represent individual reward functions, as in our model in Section 3.1. For the reward model structure of the general representation function, we froze the first 70% of the language model’s layers, using the outputs of these layers as the representation. For the linear representation function, we froze the entire language model and only trained the additional final layer. We used two clusters for the personalized reward model with user clustering. We evaluated the reward models based on their accuracy in correctly assigning higher rewards to chosen summaries over rejected summaries in the validation set. Our results, shown in Figure 2, indicate that clustering methods can efficiently learn the personalized reward model. Furthermore, personalization with general representation learning is necessary, as indicated by the performance gap compared to personalization with linear representation learning. Notably, for LLaMA3 8B, the performance differences between the Naive method and both PG and CG are statistically significant by t-test ($p<0.006$).
6.2 Experiment 2: Output Examples of Reward Aggregation with axes Dataset
In Experiment 2, we aggregated three axes rewards using Equation 4.3 with $\alpha=-\infty,-1,0,1,\infty$. We included representative outputs from these aggregated results in Section I.1.